OpenStudy (anonymous):

An isosceles triangle has base angle 24 degrees and height 19. Find the area.

OpenStudy (experimentx):

19*19*cot(24)

OpenStudy (anonymous):

Let 2x = base of triangle tan 24 = 19/ x -> x = 19/tan24 = 42.67 => A = b *h = 2* 42.67 * 19 = 1,621.5